OpenStudy (rebeccapink6398):

Which mixed number is equivalent to this decimal? 7.59

OpenStudy (rebeccapink6398):

@Destinymasha PLZ HELP :D UR NEVERRRRRR WRONG :)

OpenStudy (rebeccapink6398):

or anyone plzz

OpenStudy (jdoe0001):

\(\bf 7.59\implies \cfrac{759}{100}\)

OpenStudy (destinymasha):

-yea i am lol

OpenStudy (jdoe0001):

2 decimals, 2 zeros for the denominator

OpenStudy (rebeccapink6398):

that's not a possible answrr

OpenStudy (rebeccapink6398):

lol yeah @Destinymasha

OpenStudy (destinymasha):

do you have options?

OpenStudy (jdoe0001):

is not because you have to make it a mixed fraction

OpenStudy (jdoe0001):

notice is an "improper" so-called fraction

OpenStudy (rebeccapink6398):

yes wait one min

OpenStudy (rebeccapink6398):

^^^^THERE^^^^

OpenStudy (jdoe0001):

\(\bf 7.59\implies \cfrac{759}{100}\implies \cfrac{700}{100}+\cfrac{59}{100}\) what do you think it might be?

OpenStudy (rebeccapink6398):

yes but that's not one of the options

OpenStudy (jdoe0001):

well what do you think about \(\bf \cfrac{700}{100}\) can it be simplified?

OpenStudy (rebeccapink6398):

I don't know how to sim.

OpenStudy (jdoe0001):

100 + 100 + 100 + 100 + 100 + 100 + 100 =700 so there are 7* 100 = 700 thus \(\bf 7.59\implies \cfrac{759}{100}\implies \cfrac{700}{100}+\cfrac{59}{100}\implies \cfrac{7\cdot \cancel{100}}{\cancel{100}}+\cfrac{59}{100}\\ \quad \\ 7\cfrac{59}{100}\)

OpenStudy (rebeccapink6398):

thank you :)

OpenStudy (rebeccapink6398):

could you also help me with this one? it says to shows the numbers in order from least to greatest?

OpenStudy (rebeccapink6398):

@jdoe0001 could u plz help me?

OpenStudy (jdoe0001):

\(\bf 7\cfrac{4}{5}\implies \cfrac{7\cdot 5+4}{5}\implies \cfrac{39}{5}\implies 39\div 5\) what would that give?

OpenStudy (rebeccapink6398):

7.8?

OpenStudy (jdoe0001):

so which one do you think is biggest? 7.8, 7.6, 7.41? notice that 7.80, 7.60, 7.41

OpenStudy (rebeccapink6398):

39 divided by 5= 7.8

OpenStudy (jdoe0001):

so.... 7.80, 7.60, 7.41 <--- there, just sort them from smaller to biggest

OpenStudy (rebeccapink6398):

ok so 7.4,7.60,7.80 and last question :) Which mixed number is greater than 8.45?

OpenStudy (jdoe0001):

got any choices?

OpenStudy (rebeccapink6398):

8 3/8 8 4/9 8 3/5 8 5/12

OpenStudy (rebeccapink6398):

@jdoe0001 ....

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 8\frac{3}{8}\implies & 8+(3\div8)\\ \quad \\ 8\frac{4}{9}\implies & 8+(4\div9)\\ \quad \\ 8\frac{3}{5}\implies & 8+(3\div5)\\ \quad \\ 8\frac{5}{12}\implies & 8+(5\div12) \end{array}\) so which one do you think would be greater than 8.45?

OpenStudy (rebeccapink6398):

I don't get it..

OpenStudy (jdoe0001):

the fraction can be written the same way if you divide them like so and add them you'd end up with a decimal amount which you can compare against 8.45

OpenStudy (jdoe0001):

for example \(\large \begin{array}{llll} 8\frac{3}{8}\implies & 8+(3\div8)\implies 8+(0.375)\implies 8.375\\ \quad \\ 8\frac{4}{9}\implies & 8+(4\div9)\\ \quad \\ 8\frac{3}{5}\implies & 8+(3\div5)\\ \quad \\ 8\frac{5}{12}\implies & 8+(5\div12) \end{array}\) now, 8.375 is not greater than 8.45, so is not that one

OpenStudy (rebeccapink6398):

ok so 8 5/12 is greater?

OpenStudy (rebeccapink6398):

YAY!!! I GOT THE TEXT CORRECT THXXXXXXXX